A new cruise ship line has just launched 3 new ships: the Pacific Paradise, The caribbean paradise, and the mediterranean paradise. The carribean paradise has 32 more deluxe sytateroooms than the pacific paradise. The Mediterranean paradise has 37 few deluxe staterooms than three times the number of deluxe staterooms on the pacific paradise.Find the number of deluxe staterooms for each of the ships if the total number of deluxe staterooms for the three ships is 705
1
Expert's answer
2011-07-20T12:27:27-0400
Let's denote the number of deluxe staterooms at Pacific Paradise as X, then the number of deluxe staterooms at Caribean Paradise is X + 32, at Mediterranean Paradise is 3X-37. As the total number of deluxe staterooms is 705,
X + X + 32 +3X - 37 = 705
5X - 5 = 705
5X = 710
X = 142

Answer: the Pacific Paradise has 142 deluxe staterooms, The Caribbean Paradise has 142 + 32 = 174 deluxe staterooms, The Mediterranean Paradise has 389 deluxe staterooms.